# Fractional diffusion as the limit of a short range potential Rayleigh gas

### v1

@@ 4.1. Energy estimates
$\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|B(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x$
$=\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|B(v^{\prime}-\bar{v}^{\prime},\omega)\,\mathrm{d}\bar{v}^{\prime}\,\mathrm{d}{\omega}\,\mathrm{d}v^{\prime}\,\mathrm{d}x$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\big{[}|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}$
$\qquad\qquad\Theta(\sin\Theta)^{-1}\,\mathrm{d}\bar{v}^{\prime}\,\mathrm{d}{\omega}\,\mathrm{d}v^{\prime}\,\mathrm{d}x$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\big{[}|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}$
$\qquad\qquad\int_{\mathbb{S}^{1}}\int_{0}^{\tilde{\sigma}}\Theta\,\mathrm{d}\Theta\,\mathrm{d}\psi\,\mathrm{d}\bar{v}^{\prime}\,\mathrm{d}v^{\prime}\,\mathrm{d}x,$
where in the last equality we used the fact that $\omega$ is the direction of the apse line $\omega=(\Theta,\psi)\in[0,\frac{\pi}{2}]\times[0,2\pi]$ and $\mathrm{d}\omega=(\sin\Theta)^{d-2}\mathrm{d}\Theta\mathrm{d}\psi$, where $d$ is the dimension, i.e. here $d=3.$ Therefore, the derivative of the momentum energy is estimated by
$\frac{d}{dt}D_{f}(t)\leq\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\pi\tilde{\sigma}^{2}c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\big{[}|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}^{\prime}\,\mathrm{d}v^{\prime}\,\mathrm{d}x.$
We now start using $2<n\leq 5$. Then
$\frac{\mathrm{d}}{\mathrm{d}t}D_{f}(t)$
$\leq\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\pi\tilde{\sigma}^{2}c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\big{[}|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}^{\prime}\,\mathrm{d}v^{\prime}\,\mathrm{d}x$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\pi\tilde{\sigma}^{2}c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})\bigg{(}\int_{\mathds{R}^{3}\setminus B(v^{\prime},1)}g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\big{[}|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}\mathrm{d}\bar{v}^{\prime}$
$\quad+\int_{B(v^{\prime},1)}g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\big{[}|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}^{\prime}\bigg{)}\,\mathrm{d}v^{\prime}\,\mathrm{d}x,$
where
$\int_{\mathds{R}^{3}\setminus B(v^{\prime},1)}g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\big{[}|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}^{\prime}$ $\leq(1+\mathcal{O}(1))\int_{\mathds{R}^{3}}g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\,\mathrm{d}\bar{v}^{\prime}$
$=\pi C_{1}\bar{R}^{4}+\frac{4\pi C_{2}}{q-4}\bar{R}^{4-q}+\mathcal{O}(1),$
when $q>4$, since $|v^{\prime}-\bar{v}^{\prime}|\geq 1$ on the set $\mathds{R}^{3}\setminus B(v^{\prime},1)$ and $\frac{n-5}{n-1}$, $\frac{n-7}{n-1}\leq 0$, then $|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}$, $|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}}\leq 1$.
The other term can be estimated using $n>2$ by
$\int_{B(v^{\prime},1)}g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\big{[}|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}$ $+\mathcal{O}(|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}^{\prime}$
$=\int_{B(0,1)}g_{0}(\bar{v}^{\prime}-v^{\prime})|\bar{v}^{\prime}-v^{\prime}|\big{[}|\bar{v}^{\prime}|^{\frac{n-5}{n-1}}+\mathcal{O}(|\bar{v}^{\prime}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}^{\prime}$
$\leq 4\pi\|g_{0}(v^{\prime})(1+|v^{\prime}|)\|_{L^{\infty}((\mathds{R}^{3})}\int_{0}^{1}r^{\frac{n-5}{n-1}+2}+\mathcal{O}(1)r^{\frac{n-7}{n-1}+2}\,\mathrm{d}r$
$=\bigg{(}\frac{n-1}{n-2}+\frac{2(n-1)}{2n-5}\mathcal{O}(1)\bigg{)}\pi\|g_{0}(v^{\prime})(1+|v^{\prime}|)\|_{L^{\infty}(\mathds{R}^{3})}.$
This is bounded by the assumptions that $2<n\leq 5$ and the background distribution function $g_{0}$ has a first moment and $g_{0}(v)(1+|v|)\in L^{\infty}(\mathds{R}^{3})$. Thus,
$\frac{\mathrm{d}}{\mathrm{d}t}D_{f}(t)\leq\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\pi\tilde{\sigma}^{2}c\bigg{(}\pi C_{1}\bar{R}^{4}+\frac{4\pi C_{2}}{q-4}\bar{R}^{4-q}+\frac{n-1}{n-2}\pi\|g_{0}(v^{\prime})(1+|v^{\prime}|)\|_{L^{\infty}(\mathds{R}^{3})}+\mathcal{O}(1)\bigg{)},$
@@
by the computations above and the fact that the function $f_{t}$ is a probability density. We notice that the right hand side of the above differential inequality does not depend on $t$. Denote the right hand side by $\tilde{C}_{q,n}c$. Then, by e.g. [Tes12, Theorem 1.3] we get
@@
This means that the momentum $D_{f}(t)$ grows linearly with $t$. ∎
@@ 4.2. Number of collisions
The expected number of collisions grows linearly with $t$. That is for any $t\geq 0$ and for $2<n\leq 5$
@@
$=\int_{\mathcal{MT}}n(\Phi)c[Q_{t}^{+}[P_{t}](\Phi)-P_{t}(\Phi)Q^{-}_{\tau}(\Phi)]\,\mathrm{d}\Phi,$
[…]
$\int_{\mathcal{MT}}n(\Phi)Q_{t}^{+}[P_{t}](\Phi)\,\mathrm{d}\Phi$ $=\int_{\mathcal{MT}}n(\Phi)\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
$=\int_{\mathcal{MT}}n(\bar{\Phi})\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
$\quad+\int_{\mathcal{MT}}\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
@@
$\int_{\mathcal{MT}}n(\Phi)P_{t}(\Phi)Q^{-}_{\tau}(\Phi)\,\mathrm{d}\Phi=\int_{\mathcal{MT}}n(\Phi)P_{t}(\Phi)\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\omega\,\mathrm{d}\bar{v}\,\mathrm{d}\Phi.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}(n(\Phi))(t)$ $=c\int_{\mathcal{MT}}n(\Phi)\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
$\quad-c\int_{\mathcal{MT}}n(\Phi)P_{t}(\Phi)\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\omega\,\mathrm{d}\bar{v}\,\mathrm{d}\Phi$
$\quad+c\int_{\mathcal{MT}}\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}(n(\Phi))(t)=c\int_{\mathcal{MT}}[Q^{+}[h_{t}](\Phi)-$ $h_{t}(\Phi)Q^{-}_{\tau}(\Phi)]\,\mathrm{d}\Phi$
$+c\int_{\mathcal{MT}}\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}h_{t}(\Phi)$ $=c[Q^{+}[h_{t}](\Phi)-h_{t}(\Phi)Q^{-}_{\tau}(\Phi)]+c\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega).$
Now, let us define the forcing term $\Delta(t,\varepsilon)=\Delta^{\varepsilon}(t):=\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega).$ Then the above equation becomes
$\frac{\mathrm{d}}{\mathrm{d}t}h_{t}(\Phi)$ $=c[Q^{+}[h_{t}](\Phi)-h_{t}(\Phi)Q^{-}_{\tau}(\Phi)]+c\Delta^{\varepsilon}(t).$
@@
[…]
@@
Here, in the second equality we used Fubini’s Theorem and in the last equality that the solution semigroup $P_{t-s}$ preserves the measure. Now we will bound the term $\|\Delta^{\varepsilon}(s)\|_{L^{1}(\mathcal{MT})}$ using that $2<n\leq 5$. Then
$\|\Delta$ (${}^{\varepsilon}(s)\|_{L^{1}(\mathcal{MT})}=\int_{\mathcal{MT}}\mathds{1}_{s=\tau(\Phi)}P_{s}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
$=$ $\int_{\mathcal{MT}}\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}P_{s}(\bar{\Phi})g_{0}(\bar{v})B(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\omega\,\mathrm{d}\bar{v}\,\mathrm{d}\bar{\Phi}$
$=$ $\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\int_{\mathcal{MT}}\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}P_{s}(\bar{\Phi})g_{0}(\bar{v})\big{[}|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-7}{n-1}})\big{]}$
$\qquad\qquad\qquad\qquad\Theta(\sin\Theta)^{-1}\,\mathrm{d}\omega\,\mathrm{d}\bar{v}\,\mathrm{d}\bar{\Phi}$
$=$ $\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\pi\tilde{\sigma}^{2}\int_{\mathcal{MT}}P_{s}(\bar{\Phi})\bigg{(}\int_{\mathds{R}^{3}\setminus B(v^{\varepsilon}(\tau^{-}),1)}g_{0}(\bar{v})\big{[}|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-7}{n-1}})\big{]}\mathrm{d}\bar{v}$
$\qquad\qquad+\int_{B(v^{\varepsilon}(\tau^{-}),1)}g_{0}(\bar{v})\big{[}|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}\bigg{)}\,\mathrm{d}\bar{\Phi},$
where
$\int_{\mathds{R}^{3}\setminus B(v^{\varepsilon}(\tau^{-}),1)}g_{0}(\bar{v})\big{[}|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}$ $\leq(1+\mathcal{O}(1))\int_{\mathds{R}^{3}}g_{0}(\bar{v})\,\mathrm{d}\bar{v}$
$=\mathcal{O}(1),$
since $|v^{\prime}-\bar{v}^{\prime}|\geq 1$ on the set $\mathds{R}^{3}\setminus B(v^{\prime},1)$ and $\frac{n-5}{n-1}$, $\frac{n-7}{n-1}\leq 0$, then $|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-5}{n-1}}$, $|v^{\prime}-\bar{v}^{\prime}|^{\frac{n-7}{n-1}}\leq 1$. The other term can be estimated by
$\int_{B(v^{\varepsilon}(\tau^{-}),1)}g_{0}(\bar{v})\big{[}|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v^{\varepsilon}(\tau^{-})-\bar{v}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}$
$=\int_{B(0,1)}g_{0}(\bar{v}-v^{\varepsilon}(\tau^{-}))\big{[}|\bar{v}|^{\frac{n-5}{n-1}}+\mathcal{O}(|\bar{v}|^{\frac{n-7}{n-1}})\big{]}\,\mathrm{d}\bar{v}$
$\leq 4\pi\|g_{0}(\bar{v})\|_{L^{\infty}((\mathds{R}^{3})}\int_{0}^{1}r^{\frac{n-5}{n-1}+2}+\mathcal{O}(1)r^{\frac{n-7}{n-1}+2}\,\mathrm{d}r$
$=\bigg{(}\frac{n-1}{n-2}+\frac{2(n-1)}{2n-5}\mathcal{O}(1)\bigg{)}\pi\|g_{0}(\bar{v})\|_{L^{\infty}(\mathds{R}^{3})}.$
Therefore, as $2<n\leq 5$, the term is bounded as long as the distribution function $g_{0}$ is in $L^{\infty}(\mathds{R}^{3})$. Thus
$\|\Delta^{\varepsilon}(s)\|_{L^{1}(\mathcal{MT})}$ $\leq\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\pi\tilde{\sigma}^{2}\bigg{(}\frac{n-1}{n-2}\pi\|g_{0}(\bar{v})\|_{L^{\infty}(\mathds{R}^{3})}+\mathcal{O}(1)\bigg{)}\int_{\mathds{R}^{3}\times\mathds{R}^{3}}f_{s}(x,v)\,\mathrm{d}v\,\mathrm{d}x,$
where the integral on the right hand side is equal to $1$, since the function $f_{s}$ is a probability distribution function. We denote the right hand side term by $C$. Then, we get
@@
$\mathbb{E}(n(\Phi))(t)$ $\leq 1+Cct,\quad\textrm{for \ all \ }t\geq 0\ \textrm{and \ for\ }2<n\leq 5.$

### v2

@@ 4.1. Energy estimates
$\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\mathcal{B}(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x$
$=\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\mathcal{B}(v^{\prime}-\bar{v}^{\prime},\omega)\,\mathrm{d}\bar{v}^{\prime}\,\mathrm{d}{\omega}\,\mathrm{d}v^{\prime}\,\mathrm{d}x.$
Using the fact that the function $f_{t}$ is a probability density and that $\mathcal{B}$ is uniformly bounded by Lemma 3.2 and Proposition 3.3 and using $g_{0}(.)(1+|.|)\in L^{1}(\mathds{R}^{3})$ we have that the integral can be bounded by a constant $\tilde{C}_{q,n}$ only depending on $g_{0}$ and $n$. Thus,
$\frac{\mathrm{d}}{\mathrm{d}t}D_{f}(t)\leq c\tilde{C}_{q,n}.$
@@
We notice that the right hand side of the above differential inequality does not depend on $t$. Then, by e.g. [Tes12, Theorem 1.3] we get
@@
This means that the momentum $D_{f}(t)$ grows linearly with $t$.
@@ 4.2. Number of collisions
The expected number of collisions grows linearly with $t$. That is for any $t\geq 0$ and for $3<n\leq 5$
@@
$=\int_{\mathcal{MT}}n(\Phi)c[Q_{t}^{+}[P_{t}](\Phi)-Q^{-}_{t}[P_{t}](\Phi)]\,\mathrm{d}\Phi,$
here we used equation (2.1) and the fact that $\frac{\mathrm{d}}{\mathrm{d}t}n(\Phi)=0.$ Now,
$\int_{\mathcal{MT}}n(\Phi)Q_{t}^{+}[P_{t}](\Phi)\,\mathrm{d}\Phi$ $=\int_{\mathcal{MT}}n(\Phi)\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
$=\int_{\mathcal{MT}}n(\bar{\Phi})\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
$\quad+\int_{\mathcal{MT}}\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
@@
$\int_{\mathcal{MT}}n(\Phi)Q^{-}_{t}[P_{t}](\Phi)\,\mathrm{d}\Phi=\int_{\mathcal{MT}}n(\Phi)P_{t}(\Phi)\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau)-\bar{v},\omega)\,\mathrm{d}\omega\,\mathrm{d}\bar{v}\,\mathrm{d}\Phi.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}(n(\Phi))(t)$ $=c\int_{\mathcal{MT}}n(\Phi)\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
$\quad-c\int_{\mathcal{MT}}n(\Phi)P_{t}(\Phi)\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau)-\bar{v},\omega)\,\mathrm{d}\omega\,\mathrm{d}\bar{v}\,\mathrm{d}\Phi$
$\quad+c\int_{\mathcal{MT}}\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}(n(\Phi))(t)=c\int_{\mathcal{MT}}[Q^{+}[h_{t}](\Phi)$- $Q^{-}_{t}[h_{t}](\Phi)]\,\mathrm{d}\Phi$
$+c\int_{\mathcal{MT}}\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}h_{t}(\Phi)$ $=c[Q^{+}[h_{t}](\Phi)-Q^{-}_{t}[h_{t}](\Phi)]+c\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega).$
Now, let us define the forcing term $\Delta^{\varepsilon}(t):=\mathds{1}_{t=\tau(\Phi)}P_{t}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega).$ Then the above equation becomes
$\frac{\mathrm{d}}{\mathrm{d}t}h_{t}(\Phi)$ $=c[Q^{+}[h_{t}](\Phi)-Q^{-}_{t}[h_{t}](\Phi)]+c\Delta^{\varepsilon}(t).$
@@
where $P_{t}$ is the solution semigroup of equation (2.1). Integrating over $\mathcal{MT}$ we obtain
@@
Here, in the second equality we used Fubini’s Theorem and in the last equality that the solution semigroup $P_{t-s}$ preserves the measure. Now we will bound the term $\|\Delta^{\varepsilon}(s)\|_{L^{1}(\mathcal{MT})}$, then
$\|\Delta$ $^{\varepsilon}(s)\|_{L^{1}(\mathcal{MT})}=\int_{\mathcal{MT}}\mathds{1}_{s=\tau(\Phi)}P_{s}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\Phi$
$=$ $\int_{\mathcal{MT}}\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}P_{s}(\bar{\Phi})g_{0}(\bar{v})\mathcal{B}(v^{\varepsilon}(\tau^{-})-\bar{v},\omega)\,\mathrm{d}\omega\,\mathrm{d}\bar{v}\,\mathrm{d}\bar{\Phi}\leq C$
using the uniform boundedness of $\mathcal{B}$ from Lemma 3.2 and Proposition 3.3 as well as that $P_{s}$ is a probability measure and that $g_{0}\in L^{1}(\mathds{R}^{3})$.
Then, we get
@@
$\mathbb{E}(n(\Phi))(t)$ $\leq 1+Cct,\quad\textrm{for \ all \ }t\geq 0\ \textrm{and \ for\ }3<n\leq 5.$